The equilibrium constant, K, of a certain first If you were going to graphically determine the enthalpy, AH°, for this reaction, what points would you plot? order reaction was measured at two temperatures, T. The data is shown in this table Number Number T(K)K 3253.10 725 8.77 To avoid rounding errors, use three significant figures in the x values and four significant figures irn the y values point 1 Number Number point 2 Determine the rise, run, and slope of the line formed by these points rise run slope Number Number Number What is the standard enthalpy of this reaction? Number J/ molThe equilibrium constant, K, of a certain first order reaction was measured at two temperatures, T. The data is shown in this table.
Answer #1

as we know

\Delta H = -RT lnK

we will plot this on curve as

lnK on Y- axis & 1/T on X-axis

lnK=-\frac{\Delta H^o}{RT}

so slope = -\frac{\Delta H^o}{R}

we have

T(K) K
325 3.10
725 8.77
1/T * 10-4 lnK
30.76 1.13
13.79 2.17

therefore

Rise = Y2-Y1=2.17-1.13 =1.04

Run =X2X1=(13.79-30.76)*10-4= -0.001697

slope = rise/run =-612.84

slope *(-R) =H^o

Ho=8.134*612.84 =4984.84056 J/mol